For what values of 𝑐 does the curve 𝑓(π‘₯) = 2π‘₯^3 + 𝑐π‘₯^2 + 2π‘₯ have maximum and minimum points?




Show that the minimum and maximum points of every curve in the family of polynomials


𝑓(π‘₯) = 2π‘₯^3 + 𝑐π‘₯^2 + 2π‘₯ lie on the curve 𝑦 = π‘₯ βˆ’ π‘₯^3.




1
Expert's answer
2022-05-09T18:45:56-0400

We point out that the function increases for large values of "|x|". It is due to the fact that the term "2x^3" increases as "x" increases. Consider the derivative: "f'=6x^2+2cx+2". In case we consider the equation "f'=0", we get the following discriminant: "D=4c^2-48=4(c^2-12)". Thus, for "|c|\leq\sqrt{12}", "f'\geq0". It means, that for those values of "c" the function "f(x)" increases for all values of "x". Thus, minimum and maximum points are absent. For "|c|>\sqrt{12}" we get: "x_1=\frac{-2c-\sqrt{D}}{12}", "x_2=\frac{-2c+\sqrt{D}}{12}". The function "f" increases for "x\in(-\infty,x_1)" and for "(x_2.+\infty)". It decreases for "x\in(x_1,x_2)". Thus, "x_1" is a local maximum and "x_2" is a local minimum. Consider the points of intersection of the functions "f(x)=2x^3+cx^2+2x", "y=x-x^3". We receive: "3x^3+cx^2+x=0". As we can see, the equation is the same as the equation "f'=0". It means that minimum and maximum points can be found from intersection of curves "f(x)=2x^3+cx^2+2x" and "y=x-x^3".

Answer: In case "|c|\leq\sqrt{12}", function "f(x)" has no minimum and maximum points. For "|c|>\sqrt{12}", "x_1=\frac{-2c-\sqrt{D}}{12}", where "D=4c^2-48=4(c^2-12)", is a local maximum and "x_2=\frac{-2c+\sqrt{D}}{12}" is a local minimum.